OpenStudy (anonymous):

2y=5x+11 Find Three Possible OutComes To Graph

OpenStudy (anonymous):

2y=5x+11 y = 5/2 +11/2 give it a try and i will help you more

OpenStudy (anonymous):

*5/2x

OpenStudy (anonymous):

*5x/2

OpenStudy (mayankdevnani):

first put the value of x=0

OpenStudy (anonymous):

y = 5/2 x +11/2*

OpenStudy (mayankdevnani):

2y=5*0+11

OpenStudy (mayankdevnani):

then find the value of y

OpenStudy (mayankdevnani):

can you @DamnImBad

OpenStudy (anonymous):

Thanks ,

OpenStudy (mayankdevnani):

who

OpenStudy (anonymous):

Everyone

OpenStudy (mayankdevnani):

ok welcome